Sports shirts are frequently classified as S, M, L, XL for small, medium, large
and extra-large neck sizes. S fits a neck circumference of less than 37cm, M
fits between 37 and 40.5cm and L fits between 40.5cm and 44cm while XL fits
necks over 44cm in circumference. The neck circumference of adult males has
a normal distribution with
1
Expert's answer
2021-08-24T06:52:09-0400

If we were given mean as 40 and standard deviation as 2 and asked the proportion of shirts manufactured in each category, the following solution will apply

Small

"0<X<37"

Since "\mu = 40" and "\sigma = 2"

"= P(0<X<37)\\
= P(0-40<X-\mu<37-40)\\
= P(\frac{0-40}{2}<\frac{X-\mu}{\sigma}<\frac{37-40}{2})\\
= P(-20<Z<-1.5)\\
=0.0668\\
\implies P_{small}= 0.0668*100=7 \%"


Medium

"37<X<40.5"

Since "\mu = 40" and "\sigma = 2"

"= P(37<X<40.5)\\
= P(37-40<X-\mu<40.5-40)\\
= P(\frac{37-40}{2}<\frac{X-\mu}{\sigma}<\frac{40.5-40}{2})\\
= P(-1.5<Z<0.25)\\
=0.5319\\
\implies P_{medium}= 0.5319*100=53 \%"


Large

"40.5<X<44"

Since "\mu = 40" and "\sigma = 2"

"= P(40.5<X<44)\\
= P(40.5-40<X-\mu<44-40)\\
= P(\frac{40.5-40}{2}<\frac{X-\mu}{\sigma}<\frac{44-40}{2})\\
= P(0.25<Z<2)\\
=0.3785\\
\implies P_{large}= 0.3785*100=38 \%"


Extra-Large

"X>44"

Since "\mu = 40" and "\sigma = 2"

"= P(X>44)\\
= P(X-\mu>44-40)\\
= P(\frac{X-\mu}{\sigma}>\frac{44-40}{2})\\
= P(Z>2)\\
=0.0228\\
\implies P_{extra-large}= 0.0228*100=2 \%"